OpenStudy (anonymous):

2^1/3 divided by 54^1\3 some one help pleaseee !!!!!

OpenStudy (anonymous):

\[\frac{ 2^{\frac{ 1 }{ 3 }} }{ 54^{\frac{ 1 }{ 3 }}}\] Is this how it was written?

OpenStudy (anonymous):

no instead of the fraction it uses the radical

OpenStudy (anonymous):